# Pinpointing Important Genetic Variants via A Feature-level Filter and Group Knockoffs

### v2

@@ 1.3 Our Contribution
In this article, to address the limitations of existing knockoff filters in discovering important variants of AD in the analysis of the EADB-UKBB dataset, we propose a new knockoff filter that is more powerful in inferring conditional independence hypotheses and identifying important features from a set of strongly correlated features. Our contribution in this article is two-folded.
  • Presenting a family of conditional independence hypotheses that allow identifying important features using group knockoffs. Compared to existing conditional independence hypotheses, the new hypotheses allow the use of group knockoffs construction mechanism (Dai and Barber,, 2016) to provide informative discoveries at the level of features with higher power.
  • Developing the new knockoff filter with FDR control, high empirical power and informativeness [^2] In this article, informativeness refers to the precision of catching sets in identifying conditionally independent signals and is quantified by the size and purity of catching sets.. Specifically, we thoroughly investigate the theoretical properties of the developed knockoff filter and empirically evaluate their performance via simulated experiments.
The rest of this article is organized as follows. In Section 2, we develop our knockoff filter and perform theoretical analysis. Section 3 illustrates the empirical performance of our method and its advantage over existing alternatives in power and informativeness via simulated experiments. Finally, we apply our method to the AD analysis of the EADB-UKBB dataset described in Section 4 and demonstrate that the method exhibits high power and precision. Section 5 concludes with discussions.
@@ 2.1 Problem Statement
Consider independent and identically distributed (i.i.d.) samples $\{(\textbf{x}_{i},y_{i})|i=1,\ldots,n\}$ from a joint distribution $f(\mathbf{X},Y)$, where features $\textbf{X}=(X_{1},\ldots,X_{p})^{\mathsf{T}}$ are partitioned into $K$ disjoint groups $B_{1},\ldots,B_{K}$. Let $\mathbb{X}=[\textbf{x}_{1},\ldots,\textbf{x}_{n}]^{\mathsf{T}}$ denote the data matrix and $\textbf{y}=(y_{1},\ldots,y_{n})^{\mathsf{T}}$ be the vector of responses. Our target is to perform multiple testing on conditional independence hypotheses,
@@
where $\textbf{X}_{-B_{k}}=(\textbf{X}^{\mathsf{T}}_{B_{1}},\ldots,\textbf{X}^{\mathsf{T}}_{B_{k-1}},\textbf{X}^{\mathsf{T}}_{B_{k+1}},\ldots,\textbf{X}^{\mathsf{T}}_{B_{K}})^{\mathsf{T}}$ is the subvector of X created by excluding all features in the same group $B_{k}$ of $X_{j}$. Here, we refer $H_{j}^{(\text{fg})}$ as the conditional independence hypothesis at the feature-versus-group level as it depicts the independence between a feature and the response conditional on other feature groups which omits information of other features in the same group. Under the belief that the response $Y$ only depends on a relatively small number of features, we would like to find as many important features without making too many false discoveries with respect to $H_{j}^{(\text{fg})}$’s. In other words, our target is to obtain a rejection set $\mathcal{R}^{(\text{fg})}=\{j|H_{j}^{(\text{fg})}\text{ is rejected}\}$ such that the false discovery rate (FDR, the expectation of false discovery proportion),
$\text{FDR}^{(\text{fg})}=\mathbf{E}\{\text{FDP}^{(\text{fg})}\},\quad\text{where }\text{FDP}^{(\text{fg})}=\frac{\#(\mathcal{R}^{(\text{fg})}\cap\mathcal{H}_{0}^{(\text{fg})})}{1\vee\#\mathcal{R}^{(\text{fg})}},\quad\mathcal{H}_{0}^{(\text{fg})}=\{j|H_{j}^{(\text{fg})}\text{ is true}\},$
is controlled under the target level $\alpha$ ($10<\alpha<$ 1).
Under the same setup of Candès et al., (2018) that the distribution of features X is completely known while no information is provided regarding the conditional distribution $Y|\textbf{X}$, our method differs in the hypotheses to test (3) from existing approaches, which test either conditional independence hypotheses (1) at the feature-versus-feature level or conditional independence hypotheses (2) at the group-versus-group level. Such a difference is the origin of our method’s advantages over the existing knockoff filters. On one hand, unlike the group knockoff filter (Dai and Barber,, 2016), the conclusion provided by our method is at the feature level and is thus more informative on which features contribute more to the variation of $Y$. On the other hand, compared to those that infer $H^{(\text{ff})}_{j}$’s, our method can also utilize group knockoffs that are more liberal in construction to gain power in selecting important features. However, such advantages do not come free as new challenge is posted that no existing filters can provide control on $\text{FDR}^{(\text{fg})}$, which is elaborated and addressed in the following sections.
@@ 2.2 Feature Importance Scores Under Group Knockoffs
    (Knockoffs Construction): Construct the knockoff copy $\widetilde{\textbf{X}}$ of X such that $\perp Y|\textbf{X}$ X and the joint distribution of $(\textbf{X},\widetilde{\textbf{X}})$ is exchangeable at the group level (Dai and Barber,, 2016),
@@
    where $(\textbf{X},\widetilde{\textbf{X}})_{\text{swap}(\mathcal{S})}$ is obtained by swapping $\textbf{X}_{B_{k}}$ and $\widetilde{\textbf{X}}_{B_{k}}$ in $(\textbf{X},\widetilde{\textbf{X}})$ for all $k\in\mathcal{S}$.
@@
However, directly inputting $W_{1},\ldots,W_{p}$ to the model-X knockoff filter at feature level (Candès et al.,, 2018) cannot provide control on $\text{FDR}^{(\text{fg})}$ of the rejection set. The main reason is that under the group knockoffs model $f(\textbf{X},\widetilde{\textbf{X}})$, $W_{1},\ldots,W_{p}$ do not possess the coin-flipping property (i.e., conditional on $|W_{j}|$’s and $\text{sign}(W_{j})$’s for those false $H^{(\text{fg})}_{j}$’s, $\text{sign}(W_{j})$’s for those true $H^{(\text{fg})}_{j}$’s are independently and uniformly distributed on {±$\{\pm 1\}$), the key property that guarantees FDR control in Candès et al., (2018). Specifically, for most of the widely-used feature importance scores (e.g., the marginal correlations, the Lasso coefficient-difference statistics and the Lasso signed max statistics; Barber and Candès,, 2015, Candès et al.,, 2018), $W_{j}$’s for those true $H^{(\text{fg})}_{j}$’s are dependent coin flips. This comes from (4) where exchangeability only exists among groups, making signs of $W_{j}$’s from the same group not independent.
Thus, we relax the unachievable coin-flipping property to the between-group coin-flipping property which can be satisfied by feature importance scores $W_{j}$’s in the form of
$\begin{cases}T_{j}=t([\mathbb{X}_{j},\mathbb{X}_{B_{k}\setminus\{j\}},\widetilde{\mathbb{X}}_{j},\widetilde{\mathbb{X}}_{B_{k}\setminus\{j\}},\{\mathbb{X}_{-B_{k}},\widetilde{\mathbb{X}}_{-B_{k}}\}],\textbf{y}),\\ \widetilde{T}_{j}=t([\widetilde{\mathbb{X}}_{j},\widetilde{\mathbb{X}}_{B_{k}\setminus\{j\}},{\mathbb{X}}_{j},{\mathbb{X}}_{B_{k}\setminus\{j\}},\{\mathbb{X}_{-B_{k}},\widetilde{\mathbb{X}}_{-B_{k}}\}],\textbf{y}),\end{cases}\quad k=1,\ldots,K\text{, }j\in B_{k},$ (5)
where $\text{where }\{\mathbb{X}_{-B_{k}},\widetilde{\mathbb{X}}_{-B_{k}}\}\text{ denotes the set of unordered pairs }\{\mathbb{X}_{j^{\prime}},\widetilde{\mathbb{X}}_{j^{\prime}}\}\text{ for all }j^{\prime}\notin B_{k},$
as shown in Theorem 1.
@@ 1. $\diamond$
      (Uniformity) $\text{\rm sign}(W_{j})$ is uniformly distributed on {±$\{\pm 1\}$ for all $j\in\mathcal{H}_{0}^{(\text{fg})}$;
@@ 2. $\diamond$
The proof of Theorem 1 is provided in Section A of the Appendix. Specifically, Theorem 1 relies on a condition “$H_{j}^{(\text{fg})}$ implies $H_{k}^{\text{(gg)}}$ for all $j\in B_{k}$ and $k=1,\ldots,K$”. Such a condition is generally valid in genetic data analysis, analogous to the usual expectation in fine-mapping that the association between an important variant $X_{j}$ and the response $Y$ is in the same direction as the direct effect of $X_{j}$ on $Y$.
Examples of feature importance scores in the form of (5) include:
  • $\diamond$
  • $\diamond$
    joint lasso: absolute values of lasso estimators $T_{j}=\Big{|}\widehat{\beta}_{j}\Big{|}$ and $\widetilde{T}_{j}=\Big{|}\widehat{\widetilde{\beta}}_{j}\Big{|}$ of the linear model
  • $\diamond$
    marginal correlation with lasso residual: absolute values of estimated marginal correlations $T_{j}=\Big{|}\widehat{\rho}(X_{j},\hat{e})\Big{|}$ and $\widetilde{T}_{j}=\Big{|}\widehat{\rho}(\widetilde{X}_{j},\hat{e})\Big{|}$ between $(X_{j},\widetilde{X}_{j})^{\mathsf{T}}$ and the lasso residual of the linear model
  • $\diamond$
Although the above feature importance scores can be used to compute feature statistics $W_{1},\ldots,W_{j}$ with the between-group coin-flipping property, they suffer several problems in practice. On one hand, with not much information of $H_{j}^{(\text{fg})}$’s, marginal correlation with response usually lacks power. So do the joint lasso feature importance scores, which do not directly depict evidence against $H_{j}^{(\text{fg})}$’s. On the other hand, with the need to perform $p$ and $K$ lasso linear model fittings, separate lasso and marginal correlation with lasso residual are both computationally intensive.
Faced with such a tradeoff between computational efficiency and power, we propose the combined feature importance scores as computationally efficient proxies of the separate lasso importance scores
$\begin{cases}T_{j}=|g_{j}|\cdot|\widehat{\rho}(X_{j},Y)|,\\ \widetilde{T}_{j}=|g_{j}|\cdot|\widehat{\rho}(\widetilde{{X}}_{j},Y)|,\\ \end{cases}\quad k=1,\ldots,K\text{, }j\in B_{k},$ (7)
where $g_{j}$ is computed as follows.
  1. 1.
    Fit the joint linear model (6) and compute $\widehat{\gamma}_{j}=(\widehat{\beta}_{j}+\widehat{\widetilde{\beta}}_{j})/2$ and $\widehat{\delta}_{j}=(\widehat{\beta}_{j}-\widehat{\widetilde{\beta}}_{j})/2$ as the approximate estimator of the reparametrized model
    $Y=\sum_{j=1}^{p}{\gamma}_{j}\left({{X}}_{j}+\widetilde{{X}}_{j}\right)+\sum_{j=1}^{p}{\delta}_{j}\left({{X}}_{j}-\widetilde{{X}}_{j}\right)+e.$
  2. 2.
    For $k=1,\ldots,K$ and $j\in B_{k}$, compute
    $g_{j}$ $=\widehat{\gamma}_{j}+\sum_{j^{\dagger}\in B_{k}\setminus\{j\}}\frac{\psi_{j,j^{\dagger}}-\boldsymbol{\Psi}_{j,-B_{k}}\boldsymbol{\Psi}^{-1}_{-B_{k},-B_{k}}\boldsymbol{\Psi}_{-B_{k},j^{\dagger}}}{\psi_{j,j}-\boldsymbol{\Psi}_{j,-B_{k}}\boldsymbol{\Psi}^{-1}_{-B_{k},-B_{k}}\boldsymbol{\Psi}_{-B_{k},j}}\widehat{\gamma}_{j^{\dagger}},$
    where $\boldsymbol{\Psi}=(\psi_{j,j^{\dagger}})_{p\times p}=\boldsymbol{\Sigma}-\textbf{D}/2,\quad\boldsymbol{\Sigma}=\text{Var}(\textbf{X}),\quad\textbf{D}=\boldsymbol{\Sigma}-\text{Cov}(\textbf{X},\widetilde{\textbf{X}}),$
    as the approximate estimator of the model
    $Y={g}_{j}\left({{X}}_{j}+\widetilde{{X}}_{j}\right)+\sum_{j^{\prime}\notin B_{k}}{\gamma}_{j^{\prime}}\left({{X}}_{j^{\prime}}+\widetilde{{X}}_{j^{\prime}}\right)+{h}_{j}\left({{X}}_{j}-\widetilde{{X}}_{j}\right)+\sum_{j^{\prime}\notin B_{k}}{\delta}_{j^{\prime}}\left({{X}}_{j^{\prime}}-\widetilde{{X}}_{j^{\prime}}\right)+e.$
    In other words, ${g}_{j}$ measures the overall contribution of ${X}_{j}$ and $\widetilde{{X}}_{j}$ to the variation of $Y$ beyond features from other groups.
By doing so, feature statistics $W_{j}=w_{j}(T_{j},\widetilde{T}_{j})$ ($j=1,\ldots,p$) with antisymmetric functions $w_{j}$’s satisfy the between-group coin-flipping property because
$\text{sign}(W_{j})=\text{sign}(w_{j}(|\widehat{\rho}({{X}}_{j},Y)|,|\widehat{\rho}(\widetilde{{X}}_{j},Y)|)),$
whose between-group coin-flipping property is validated in Theorem 1. In addition, as $g_{j}$ is derived to measure the total dependency between $(X_{j},\widetilde{{X}}_{j})$ and the response $Y$ conditional on $(\textbf{X}_{-B_{k}},\widetilde{\textbf{X}}_{-B_{k}})$, $|W_{j}|$’s computed from the combined feature importance scores provide proxies of $|W_{j}|$’s computed from the separate lasso feature importance scores.
2.3 Naive Feature Filters under Group Knockoffs
@@
By Theorem 1, feature statistics $W_{j}$’s only satisfy the between-group coin-flipping property and thus directly applying the model-X knockoff filter (Candès et al.,, 2018) would violate the FDR control. However, we can still utilize the between-group independence among $W_{j}$’s for inference with FDR control as follows.
@@
Recognizing that independence remains true among $W_{j}$’s for those true $H^{(\text{fg})}_{j}$’s from different groups, we consider aligning $W_{j}$’s in Table 2 such that $W_{j}$’s for all features in the group $B_{k}$ are in the $k$-th column and $|W_{(k1)}|\geq|W_{(k2)}|\geq\cdots$. Let $\mathcal{C}_{l}$ be the set of features whose feature statistics are aligned in the $l$-th row and consider the rejection set in the form of $\mathcal{R}^{(\text{fg})}(t)=\{j|W_{j}\geq t\}$.
@@
Row $B_{1}$ $B_{2}$ $B_{3}$cdots
1 $W_{(11)}$ $W_{(21)}$ $W_{(31)}$cdots
[…]
  • On one hand, we find that within each row, there exists at most one $W_{j}$ from each group. As feature statistics $W_{j}$’s satisfy the between-group coin-flipping property, we have that $W_{j}$’s within the same row satisfy the coin-flipping property introduced in Candès et al., (2018). Thus, we can estimate the false discovery proportion (FDP) of the rejection subset $\mathcal{R}^{(l)}(t)=\{j\in\mathcal{C}_{l}|W_{j}\geq t\}$ as
@@ 1 $W_{(11)}$ $W_{(21)}$ $W_{(31)}$cdots
    for any $t>0$ and $l=1,2\ldots$ in the analogous way of Candès et al., (2018).
@@
    where the second equality comes from the fact that rejection subsets $\mathcal{R}^{(l)}(t)$ can be nonempty only for the first $\phi(t)=\max_{k}\#\{j\in B_{k}||W_{j}|\geq t\}$ rows because $|W_{(k1)}|\geq|W_{(k2)}|\geq\cdots$ with each group $B_{k}$.
@@
2.4 Rigorous Feature Filters under Group Knockoffs
Although the naive filter seems intuitive in FDR control, such a control can not be proven. To fill the gap, we provide an alternative interpretation of Algorithm 1 and consider a variation that provides provable FDR control as follows.

### v3

@@ 1.3 Our Contribution
In this article, to address the limitations of existing knockoff methods in discovering important variants of AD in the analysis of the EADB-UKBB dataset, we propose a new knockoff filter that achieves balance between the statistical power and precision in identifying important features from a set of strongly correlated features. Built upon group knockoffs, our filter simultaneously selects groups containing important features and performs fine-mapping to identify the most promising features within each selected group. Thus, our method can return refined catching sets that are small in size and high in purity without large power loss. Specifically, our approach of refinement is motivated by penalized regressions that can efficiently learn the sparse nature of causality via the sparse feature importance, which is leveraged in searching of catching sets under alternative variant-level hypotheses.
The rest of this article is organized as follows. In Section 2, we develop our new knockoff filter and perform theoretical analysis. Section 3 illustrates the empirical performance of our method and its advantage over existing alternatives in power and precision via simulated experiments. Finally, we apply our method to the AD analysis of the EADB-UKBB dataset in Section 4 and demonstrate that the method exhibits balance between power and precision. Section 5 concludes with discussions.
@@ 2.1 Problem Statement
Consider independent and identically distributed (i.i.d.) samples $\{(\textbf{x}_{i},y_{i})|i=1,\ldots,n\}$ from a joint distribution $f(\mathbf{X},Y)$, where features $\textbf{X}=(X_{1},\ldots,X_{p})^{\mathsf{T}}$ are partitioned into $K$ disjoint groups $B_{1},\ldots,B_{K}$ for inference. For example, in analysis of genetic data, features (variants) with high correlation are usually allocated to the same group such that features from different groups are weakly or moderately correlated. Let $\mathbb{X}=[\textbf{x}_{1},\ldots,\textbf{x}_{n}]^{\mathsf{T}}$ denote the data matrix and $\textbf{y}=(y_{1},\ldots,y_{n})^{\mathsf{T}}$ be the vector of responses. Our target is to perform multiple testing on conditional independence hypotheses,
@@
where $\textbf{X}_{-B_{k}}=(\textbf{X}^{\mathsf{T}}_{B_{1}},\ldots,\textbf{X}^{\mathsf{T}}_{B_{k-1}},\textbf{X}^{\mathsf{T}}_{B_{k+1}},\ldots,\textbf{X}^{\mathsf{T}}_{B_{K}})^{\mathsf{T}}$ is the subvector of X created by excluding all features in the same group $B_{k}$ of $X_{j}$. Here, we refer $H_{j}^{(\text{fg})}$ as the conditional independence hypothesis at the feature-versus-group level as it depicts the independence between a feature and the response conditional on other feature groups which omits information of other features in the same group. Under the belief that the response $Y$ only depends on a relatively small number of features as validated in many GWAS studies, we would like to find as many important features as possible without making too many false discoveries with respect to $H_{j}^{(\text{fg})}$’s. In other words, our target is to obtain a rejection set $\mathcal{R}^{(\text{fg})}=\{j|H_{j}^{(\text{fg})}\text{ is rejected}\}$ such that the false discovery rate (FDR, the expectation of false discovery proportion),
$\text{FDR}^{(\text{fg})}=\mathbf{E}\{\text{FDP}^{(\text{fg})}\},\quad\text{where }\text{FDP}^{(\text{fg})}=\frac{\#(\mathcal{R}^{(\text{fg})}\cap\mathcal{H}_{0}^{(\text{fg})})}{1\vee\#\mathcal{R}^{(\text{fg})}},\quad\mathcal{H}_{0}^{(\text{fg})}=\{j|H_{j}^{(\text{fg})}\text{ is true}\},$ (4)
is controlled under the target level $\alpha$ ($0<\alpha<1$).
Our method is under the same setup of Candes2018 that the distribution of features X is known completely or well approximated up to the first two moments, while no information is provided regarding the conditional distribution $Y|\textbf{X}$. That is to say, our framework does not assume that the contribution of $X_{j}$’s to $Y$ are additive (independent) and the FDR control remains valid even the underlying conditional distribution $Y|\textbf{X}$ is deviated from the linear model. Our method differs from existing approaches, which test either (1) at the feature-versus-feature level or (2) at the group-versus-group level, as follows.
  • Compared to those that infer $H^{(\text{ff})}_{j}$’s, our method can utilize group knockoffs that are more liberal in construction to achieve higher power in identifying important features.
  • It is true that in most of GWAS studies where highly correlated features are allocated in the same groups, our hypotheses $H_{j}^{(\text{fg})}$’s are equivalent to $H_{k}^{(\text{gg})}$’s at the group-versus-group level for $j\in B_{k}$. That is to say, including feature $X_{j}$ in $\mathcal{R}^{(\text{fg})}$ implies that the group $B_{k}$ contains important features. The difference is that our method simultaneously selects groups containing important features and performs fine-mapping to identify the most promising features within each selected group. Thus, instead of returning variant groups with many variants as large and impure catching sets, our method can further refine the selected variant groups. In addition, our method can achieve better balance between power and precision compared to the two-step heuristic alternative – applying post-selection fine-mapping to groups selected by the group knockoff filter.
However, such a gain does not come free as new challenge is posted that no existing filters can provide control on $\text{FDR}^{(\text{fg})}$, which is elaborated and addressed in the following sections.
@@ 2.2 Feature Importance Scores Under Group Knockoffs
[…]
@@
[…]
@@
[…]
Thus, we relax the unachievable coin-flipping property to the between-group coin-flipping property and propose feature statistics $W_{j}$’s using anti-symmetric functions $w_{j}$’s and
[…]
Here, the swapping between $T_{j}$ and $\widetilde{T}_{j}$ requires the swapping between $\mathbb{X}_{B_{k}}$ and $\widetilde{\mathbb{X}}_{B_{k}}$, while it only requires the swapping between $\mathbb{X}_{j}$ and $\widetilde{\mathbb{X}}_{j}$ in the model-X knockoff filter. However, most of commonly used feature importance scores in the model-X knockoff filter also satisfy (6). Examples include:
  • $\diamond$
  • $\diamond$
    joint lasso: absolute values of lasso estimators $T_{j}=\Big|\widehat{\beta}_{j}\Big|$ and $\widetilde{T}_{j}=\Big|\widehat{\widetilde{\beta}}_{j}\Big|$ of the linear model (Tibshirani1996)
  • $\diamond$
[…]
  • $\diamond$
In our practice, we suggest using joint lasso importance score because it is computationally efficient and induces sparsity for better refinement of catching sets as stated in Section 2.1.
As shown in Theorem 1, feature importance scores in the form of (6) satisfy the between-group coin-flipping property.
@@ 1. $\diamond$
      (Uniformity) $\text{\rm sign}(W_{j})$ is uniformly distributed on $\{\pm 1\}$ for all $j\in\mathcal{H}_{0}^{(\text{fg})}$;
@@ 2. $\diamond$
The proof of Theorem 1 is provided in Appendix B. Specifically, Theorem 1 relies on a condition $H_{j}^{(\text{fg})}$ implies $H_{k}^{\text{(gg)}}$ for all $j\in B_{k}$ and $k=1,\ldots,K$”. That is to say, if $H_{k}^{\text{(gg)}}$ is false and group $B_{k}$ is important, all $H_{j}^{(\text{fg})}$’s ($j\in B_{k}$) are false and any $j\in B_{k}$ included in the selection sets would not be counted as false discovery. Such a condition is generally valid in genetic data analysis, analogous to the usual expectation in fine-mapping that the association between an important variant $X_{j}$ and the response $Y$ is in the same direction as the direct effect of $X_{j}$ on $Y$.
2.3 Naive Feature Filter under Group Knockoffs
@@
By Theorem 1, feature statistics $W_{j}$’s only satisfy the between-group coin-flipping property and thus directly applying the model-X knockoff filter (Candes2018) would violate the FDR control. However, we can still utilize the between-group independence among $W_{j}$’s for inference with FDR control as follows.
@@
Recognizing that independence remains true among $\text{sign}(W_{j})$’s for those true $H^{(\text{fg})}_{j}$’s from different groups, we consider aligning $W_{j}$’s in Table 2 such that $W_{j}$’s for all features in the group $B_{k}$ are in the $k$-th column and $|W_{(k1)}|\geq|W_{(k2)}|\geq\cdots$. Let $\mathcal{C}_{l}$ be the set of features whose feature statistics are aligned in the $l$-th row and consider the rejection set in the form of $\mathcal{R}^{(\text{fg})}(t)=\{j|W_{j}\geq t\}$.
@@
Row $B_{1}$ $B_{2}$ $B_{3}$ $\cdots$
1 $W_{(11)}$ $W_{(21)}$ $W_{(31)}$ $\cdots$
2 $W_{(12)}$ $W_{(22)}$ $W_{(32)}$ $\cdots$
3 $W_{(13)}$ $W_{(23)}$ $W_{(33)}$ $\cdots$
4 $W_{(14)}$ $W_{(24)}$ $W_{(34)}$ $\cdots$
$\vdots$ $\vdots$ $\vdots$ $\vdots$ $\ddots$
  • On one hand, we find that within each row, there exists at most one $W_{j}$ from each group. As feature statistics $W_{j}$’s satisfy the between-group coin-flipping property, we have that $W_{j}$’s within the same row satisfy the coin-flipping property introduced in Candes2018. Thus, we can estimate the false discovery proportion (FDP) of the rejection subset $\mathcal{R}^{(l)}(t)=\{j\in\mathcal{C}_{l}|W_{j}\geq t\}$ as
@@ 4 $W_{(14)}$ $W_{(24)}$ $W_{(34)}$ $\cdots$
    for any $t>0$ and $l=1,2\ldots$ in the analogous way of Candes2018.
@@
    where the second equality comes from the fact that rejection subsets $\mathcal{R}^{(l)}(t)$ can be nonempty only for the first $\phi(t)=\max_{k}\#\{j\in B_{k}||W_{j}|\geq t\}$ rows because $|W_{(k1)}|\geq|W_{(k2)}|\geq\cdots$ in each group $B_{k}$.
@@
2.4 FVG Filter
Although the naive filter seems intuitive in FDR control, such a control can not be strictly proven. To fill the gap, we provide an alternative interpretation of Algorithm 1 and consider a variation that provides provable FDR control as follows.